anna is now four times as old as her brother jong. in four years, anna's age will be twice as jong's age then. how old is each now

Answer by ewatrrr(24785)


Hi,
anna is now four times as old as her brother jong
Let x represent jong's age now and 4x Anna's age
Write as we Read***
4x + 4 = 2(x+4)
Solving for x
4x + 4 = 2x + 8
2x = 4
x = 2, Jong's age now, Anna is 8 (4*2)
checking our answer
8 + 4 = 2*6 = 12